OpenStudy (pulsified333):

The line y=−7x+0 is tangent to the graph of y=7/(x+2) at the point___?

OpenStudy (kayders1997):

@zepdrix

OpenStudy (pulsified333):

@ganeshie8

OpenStudy (kayders1997):

I'm looking in my book :P

OpenStudy (pulsified333):

:D

OpenStudy (kayders1997):

They take points away if you get help?

OpenStudy (pulsified333):

no why?

OpenStudy (pulsified333):

if i get it wrong after 10 attempts i dont get points

ganeshie8 (ganeshie8):

whats the slope of the given tangent line ?

OpenStudy (pulsified333):

-7

ganeshie8 (ganeshie8):

What do you know about the relation between `slope of tangent line` and the `derivative of the funtion` ?

OpenStudy (pulsified333):

the tangent line has the derivative in it

ganeshie8 (ganeshie8):

Here is the relation : `slope of the tangent line` at a particular point = `derivative of the function` at that point

OpenStudy (pulsified333):

oh

ganeshie8 (ganeshie8):

In other words, derivative IS the slope of tangent line

ganeshie8 (ganeshie8):

y = 7/(x+2) y' = ?

OpenStudy (pulsified333):

y'=7(-1/(x+2)^2)

ganeshie8 (ganeshie8):

Yes, set that equal to the slope of given tangent line and solve `x`

OpenStudy (pulsified333):

do i set it equal to -7x?

ganeshie8 (ganeshie8):

slope of the given tangent line is -7

OpenStudy (pulsified333):

oh okay

OpenStudy (pulsified333):

I got x=-1,-3

ganeshie8 (ganeshie8):

Yes, plug them in the function and find the corresponding y values

OpenStudy (pulsified333):

(-1,7),(-3,21)?

OpenStudy (pulsified333):

or do i plug them into the other one?

ganeshie8 (ganeshie8):

Pick the point that satisfies both

ganeshie8 (ganeshie8):

does (-1, 7) lie on y=7/(x+2) too ?

OpenStudy (pulsified333):

thank you I choose (-1,7) because it satisfied both

ganeshie8 (ganeshie8):

Yep!